# DrawTalking: Building Interactive Worlds by Sketching and Speaking

### v3

@@ 8. Conclusion
We have introduced DrawTalking, an approach to building interactive worlds by sketching and speaking, and have shown its potential. Our interface was inspired by our natural interplay between sketching, language, and our ability to communicate via make-believe. There are many exciting directions and we would be excited to see future research build on our approach or uncover other human-centered approaches to extend natural human abilities. We consider this project just one possible step and we hope that it will foster fruitful discussion and research.
@@ Appendix A Formative Steps Additional Content
Figure 14. Informal Formative Content Exploration: (top) subset of prepared and (bottom) live-performed sketch-based content for inspiration.
Figure 15. Formative Sketching Session, P1 Final Image. P1 narrated their own bird-watching story while sketching. Their iterative workflow and visual style of moving between ”islands of locations” helped inspire some features: e.g. discrete objects, object hierarchies, and spatially-flexible object relationships. Drawn roughly in-order: 1) Inline text titling the ”story”; 2) drawing of the city Toronto to set the location; 3) P1 visited the forest in Toronto; 4) P1 saw a bird, drew arrows indicating object relation (on a tree, in the forest); 5) P1 (the person sketch) found a pond and watched ducklings; 6) P1 took 200 photos.
@@ B.1. Participants’ Background
P1expl:
An experienced professor (at a higher-level institution) in computer science, interactive graphics, and new visual media (e.g. AR, VR storytelling and games) — focuses on teaching students new to programming. Uses many visual aids in classes.
P2expl:
Digital fine arts university student (design and computer science), drawing is a hobby. Uses sketching for game design; planning gameplay features and figuring out how they should work. e.g. diagramming and framework design. Experienced with tablet drawing.
P3expl:
Experience as a digital designer, economics and digital art university student, not professionally-trained in art, uses a digital canvas (Procreate) to record visual ideas.
P4expl:
Tech / Design university student. Does design for web, desktop, apps in-general. Has done digital illustration and drawings often since childhood.
P5expl:
A very experienced professor in computer programming for interactive graphics. Does live-coding and streaming for internet-based education as well. Uses and authors open-source coding libraries for interactive graphics. Teaches programming for creating visuals rather than designing visuals by hand.
P6expl:
Student in interactive media. Draws and paints using the physical medium, with lifelong interest and experience. Experienced with digital graphical design work. Low experience with programming.
P7expl:
Industry creative expert. 15+ years experience as a set designer and in the theater industry. Digital illustrator, spatial experience designer, AR/VR immersive projects, projects in entertainment and fine-arts. Creative directing, collaborations with companies and academia. Non-programmer, some experience with visual blocks-based interfaces.
P8expl:
Robotics and machine-learning researcher, works with developing ML models, often needs to visualize ML inputs/outputs and illustrates rough sketches to think-through policies (for robots). No prior experience using a tablet.
P9expl:
Robotics and machine-learning PhD student, used tablet-based sketching interfaces.
@@ B.2. Select Screenshots
Figure 17. P2: Squirrel Climbing Tree Stopped by Collider. A rule is assigned to the collider to make it a prop the user can move around themselves to stop the squirrel from climbing upon collision with the collider.(The collider is shown in the semantics diagram as a rectangle, but invisible in the canvas on-purpose — where the label ”collider” is.)
@@ Appendix C Exploratory Features and Examples for the Record
Figure 24. Text objects as macros. We can reduce repetitive speech by storing text as inline objects that work exactly as text in the semantics diagram (Figure 4). They’re created by linking the text transcript’s current content and the canvas. Here, inline text objects (in purple) have been created from 1) the word ”car” and 2) the sentence, ”the car moves right for five seconds.” The user can link a new object with 1) to label it ”car” and link 2) and the language action button to treat the text as a reusable command. (Running ”the car moves right…” again.) There is potential to explore this further. However, we decided to focus on the more general capability of triggering commands with rules, e.g. upon pressing a button.
Figure 25. Context-sensitive representations based on labels: We briefly explored using the labels to control the visual representation of number sketches. In this case, arrows connected from a freehand sketch output the sketch’s angle to the numbers, and the labels determine whether to display the angle in degrees or radians. ”Dynamic representations” was left to future work.
@@
Figure 26. Molecule Matching Lesson Prototype - This example shows a checkbox (initially empty: a) that has automatically been filled (b) after the user has recreated the structure of a sketched water ($H_{2}O$) molecule formed from sketches labeled atoms, hydrogen, and oxygen.
@@ Appendix D Abstract Semantic Structure
The following defines the basic structure of S2 (subsubsection 4.6.2), which is a simplified semantic structure graph representing an interpretable command. The concrete implementation of DrawTalking traverses this structure to generate final execution commands S3 (subsubsection 4.6.2).

### v4

@@ 9. Conclusion
We have introduced DrawTalking, an approach to building and controlling interactive worlds by sketching and speaking while telling stories, and have shown its potential. Our interface was inspired by our natural interplay between sketching and language, and our ability to communicate via make-believe. There are many possible directions, and we are excited to see future research build on our approach and uncover other human-centered approaches to extending natural human abilities. We consider this project just one possible step and we hope that it will foster fruitful discussion and research.
Acknowledgements.
We would like to thank the anonymous reviewers (of this version and an earlier CHI LBW version of this work (Rosenberg et al., 2024)) for their constructive comments; Professor Lakshminarayanan Subramanian for his feedback; Devamardeep Hayatpur for additional feedback on writing and visuals for the CHI LBW; Daniel Zhang for brainstorming and discussions; the user study participants for their time and insights; and the many others who participated in conversations and gave insights during the journey of this work.
@@ Appendix A Formative Steps Additional Content
Figure 14. Formative Content Exploration: (top) subset of prepared and (bottom) live-performed sketch-based content
Attribution and Disclaimer: All rights belong to the original creators of the referenced video content. The screen-shots are for the sole purpose of providing examples of content used as partial inspiration for our research.
a:
©Neptune Studios — minutephysics — Antimatter Explained(minutephysics, 2014)
b:
©Neptune Studios — minutephysics — Immovable Object vs. Unstoppable Force - Which Wins?(minutephysics, 2013)
c:
©Bill Wurtz — history of japan(bill wurtz, 2016)
d, e:
©Bill Wurtz — history of the entire world, i guess(bill wurtz, 2017)
f:
©Neptune Studios — minutephysics — The Physics of Windmill Design(minutephysics, 2021)
g:
©Neptune Studios — MinuteEarth — Why These Bears “Waste” Food(MinuteEarth, 2021)
h:
©Vi Hart — The Case for Hovercars(Vihart, 2016)
i:
©MIT OpenCourseWare — MIT 22.01 Introduction to Nuclear Engineering and Ionizing Radiation, Fall 2016 — 11. Radioactivity and Series Radioactive Decays(MIT OpenCourseWare, 2019b)
j:
©MIT OpenCourseWare — MIT 22.01 Introduction to Nuclear Engineering and Ionizing Radiation, Fall 2016 — 1. Radiation History to the Present — Understanding the Discovery of the Neutron(MIT OpenCourseWare, 2019a)
k:
©MIT OpenCourseWare — Atomic Theory of Matter — Lec 1 — MIT 5.111 Principles of Chemical Science, Fall 2005(MIT OpenCourseWare, 2008)
l:
©MIT OpenCourseWare — MIT RES.8-007 Cosmic Origin of the Chemical Elements, Fall 2019 — Ep. 6: Element Production (Fusion) – Part 1(MIT OpenCourseWare, 2019c)
m:
©Michael New — The Circle of Fifths - How to Actually Use It(Michael New, 2015)
n:
©Khan Academy — Organelles in eukaryotic cells — The cellular basis of life — High school biology — Khan Academy(Khan Academy, 2017)
o:
©Khan Academy — DNA — Biomolecules — MCAT — Khan Academy(Khan Academy, 2014)
[…]
[Prepared and Live-Performed Example Motivating Content]Prepared and Live-Performed Example Motivating Content: These examples are screen-shots from videos (attribution in the main figure text). Two panels of example content. Panel ”Prepared”: a) live-sketched animation demonstrating particle-antiparticle collision and annihilation; b) animation of sketches of gift boxes with red ribbons colliding into each other; c) world map of japan with cartoon images of the mongols (represented as a horseback soldier and labeled with the text ”mongols”) advances towards a figure of a Japanese solider, only to be swept-up in a tornado cartoon image; d) sequence of top-down sketch art of the solar system with the labels ”rocks,” ”ice,” and ”funny clouds.” beside the objects; e) 2D art, sideview, showing the early Earth with the sun shooting a laser at the Earth to represent how dangerous it was before the atmosphere developed. The sentence ”the sun is a deadly lazer” is written above. Another frame from the same video shows the ozone layer now developed (labeled with the text ”ozone”) between the sun and the Earth. Text is shown: ”not anymore, there’s a blanket.” (referring to the ozone layer); f) hand-drawn sketch art of a person beside a spinning windmill; g) a sequence of gulls (side-view) diving into the ocean to catch fish. All at once the steps are shown in an arrow flow diagram: gull before diving, arrow from to, gull in water getting fish, arrow from to, gull above water with fish in gull’s stomach; h) a triangular character drawn with markers on physical paper (resting above a physical notebook) encounters a car and (in a second panel) a hovercar with a smiling face - all physical paper; i) two panels of a lecturer narrating over drawing an atomic nucleus with electron shells, ”you have some parent nucleus and you’ve” (the nucleus is drawn on the blackboard) ”got a lot of electron shells” (the electron shells have been drawn around the nucleus); j) lecturer showing a formation of particles labeled n and p (two each) being brought together while speaking, ”and bring them together”; k) lecturer black-board drawing with chalk while speaking - the drawing is of a function on a graph; l) a lecturer presenting while facing the audience of the video through a transparent whiteboard with text for ”Telescopes & observing” and sketches of periodic table elements (rectangular entries) and a hill; m) music-theory-related drawing of circle of fifths on a whiteboard, with the presenter playing the notes on a physical MIDI piano keyboard on-screen; n) digital blackboard lesson, sketch of a cell. text above: ”Eukaryotic Cells - membrane bound organelles. The parts of the cell have text labels with arrows pointing to them. (e.g. DNA, nucleus). The narrator says he’s about to discuss the nucleus; o) two panels of a digital blackboard lesson on DNA structures. Screen 1 shows a sketch of 3 DNA helices. The middle one has the bases (adenine thymine, guanine cytosine) highlighted. On the left is the text: Deoxyribonucleic Acid (DNA). Below is a color key for the base pairs. Adenine (written in red) - Thymine (written in green), below, Guanine (written in blue) - Cytosine (written in yellow). The second panel shows a pre-rendered animated GIF image of a DNA structure. (The narrator says ”I found this animated gif” to introduce it. On the left, an example of DNA mappings has been drawn in the form of 2 rows of A T G C all contained in a rectangle. The columns of the rows form T,A, A,T, A,T, T,A, G,C, G,C C,G, A,T, A,T, A,T, G,C, C,G, T,A, A,T.
[…]
@@ B.1. Participants’ Backgrounds
[…]
    professionally-trained in art, uses a digital drawing application (Procreate) to record visual ideas.
[…]
@@ B.2. Select Screenshots
[…]
[p1ExploreRule1]Rule 1 is created from the user input (shown in the transcript view): ”When boys collide with trees, boys move to houses.”
@@
[…]
[p1ExploreRule2]Rule 2 is created from the user input (shown in the transcript view): ”When boys collide with houses, boys move to trees.”
@@
[…]
[Participant P1 defines infinite loop-like behavior using rules.]Participant P1 defines infinite loop-like behavior using rules. The scene contains a tree, seagulls, sun, boy, and house sketch.
Figure 17. P2: Squirrel Climbing Tree Stopped by Collider. A rule is assigned to the collider to make it a prop the user can move around themselves to stop the squirrel from climbing upon collision with the collider. (The collider is shown in the semantics diagram as a rectangle, but invisible in the canvas on-purpose — where the label ”collider” is.) P2 also created a button that causes the dog to move.
[…]
[Squirrel climbing scene]Squirrel climbing scene: canvas: sketches for a squirrel, sun, acorn, a button with the word ”chase” on it, a dog, a tree, and an invisible collider. Transcript view user input: ”when the squirrel collides with the collider the squirrel stops.”
@@
[…]
[A scene in which birds follow spawning seeds]A scene in which birds follow spawning seeds: canvas: sketch of a boy, two grass lawns, two seeds (one on the rightmost grass, a house above the rightmost grass, three birds, vertically stacked above boy on left, boy is on the left grass. Transcript view input: ”when seeds appear birds follow seeds.”
@@
[…]
[This scene shows a cat jumping on a mountain.]This scene shows a cat jumping on a mountain. The canvas consists of sketches for a cat, star, square, heart, mountain, arrow (labeled with adjective ”pointy”) (roughly left to right). Transcript view input: ”The cat jumps on the mountain two times.”
@@
[…]
[A revolving planet]A revolving planet: sketches: Saturn, Pluto, Sun, alien (very fast), moon - alien stands on the moon on right. Transcript view input: ”pluto revolves around the sun.”
@@
[…]
[A snake follows a bird]A snake follows a bird. Cartoon characters and props: sketches: a sunset backdrop, on left of vertically-stacked television, bird, snake, rocket (very fast) on right, a rock on right. Transcript view (input): ”the snake follows the bird.” The snake is selected by touch.
@@
[…]
[Outer-Space scene on the planet Mars]Outer-Space scene on the planet Mars: sketches for the surface of Mars, stars (several stars in a single sketch), asteroids, a robot, a triangle character, a duck, and an unnamed flower.
@@
[…]
[A scene with characters and props at-home/out-doors]A scene with characters and props at-home/out-doors: tree, boy, chair, table, cube, chair (comfy), girl; Transcript view (input): ”The girl jumps on the comfy chair.” The semantics diagram shows the system correctly selecting the comfy chair sketch instead of the chair that wasn’t called comfy.
@@ Appendix C Exploratory Features and Examples for the Record
Figure 24. Text objects as macros: We can reduce repetitive speech by storing text as inline objects that work exactly as text in the semantics diagram (Figure 4). They’re created by linking the text transcript’s current content and the canvas. Here, inline text objects (in purple) have been created from 1) the word ”car” and 2) the sentence, ”the car moves right for five seconds.” The user can link a new object with 1) to label it ”car” and link 2) and the language action button to treat the text as a reusable command. (Running ”the car moves right…” again.) There is potential to explore this further. However, we decided to focus on the more general capability of triggering commands with rules, e.g. upon pressing a button.
[…]
[This shows how a text sketch can be used to repeat the same command repeatedly on a car sketch by user interaction between the text sketch and the car sketch]This shows how a text sketch can be used to repeat the same command repeatedly on a car sketch by user interaction between the text sketch and the car sketch; a car sketch named ”car”, a first text sketch with the text ”car,” a second text sketch with the text ”the car moves right for five seconds.” 2 panels. panel 1: Pen+touch between the unnamed car sketch and text sketch 1 gives the unnamed car sketch the label, ”car.” panel 2: Pen+touch between text sketch 2 and the language action button in the DrawTalking interface causes DrawTalking to treat the text sketch’s raw ”the car moves right for for five seconds” text as input for a command.
Figure 25. Context-sensitive representations based on labels: We briefly explored using the labels to control the visual representation of ”number sketches.” In this case, arrows connected from a freehand sketch output the sketch’s angle to the numbers, and the labels determine whether to display the angle in degrees or radians. ”Dynamic representations” was left to future work.
[…]
[A polygon sketch connected to 2 number sketches labeled for radians and degrees respectively is shown being rotated by the user, causing the number sketches to display the shape’s angle. ]A polygon sketch connected to 2 number sketches labeled for radians and degrees respectively is shown being rotated by the user, causing the number sketches to display the shape’s angle. Arrows from the polygon to two number sketches labeled ”degree” and ”radian” indicate data flow of the numerical rotation angle of the polygon sketch to the number targets. The degree sketch shows a value of 30.415, whereas the radian sketch shows the equivalent value in radians 0.531. This shows that the number sketches interpret the same input value differently according to their names.
@@
Figure 26. Molecule Matching Lesson prototype: This example shows a checkbox (initially empty: a) that has automatically been filled (b) after the user has recreated the structure of a sketched water ($H_{2}O$) molecule formed from sketches labeled atoms, hydrogen, and oxygen.
[…]
[A molecule matching game is shown consisting of atoms and a checkbox that is marked when the user correctly recreates the user-drawn target molecule (H-2-O)]A molecule matching game is shown consisting of atoms and a checkbox that is marked when the user correctly recreates the user-drawn target molecule (H-2-O); panel a) - initial state, a box is empty (the empty checkbox) beside three sketches named hydrogen, hydrogen, oxygen. panel b) - final state, the user has named the aforementioned sketches as ”atom” and connected them into one structure. The box has transformed into a checkbox because the sketch structure matches that of water (H-2-O).
Towards additional exploration of error-tolerance, we implemented an early version of unknown verb substitution for an extension of the semantics diagram Figure 27. Unknown verbs must be substituted in the semantics diagram for the command to proceed. Once the verb is substituted, the mapping persists so as not to interrupt the user again. The idea is to interrupt the user only when necessary (i.e. there is ambiguity.) (The suggestions are found simply using WordNet (Miller, 1995) within NLTK(Loper and Bird, 2002).) Context-sensitive error recovery components like this could be explored more deeply.
Figure 27. Substitution suggestions for unknown verbs: This extension to the semantics diagram is an early exploration in error-tolerance that could be developed. When unknown verbs are used in a command, the semantics diagram displays a list of potentially-equivalent words from the existing set, for each unknown verb. The user can select one or opt to cancel. Once the verb is substituted, the mapping persists so as not to interrupt the user again. (Here, the command is ”The frog hops to a lily”, but ”hop” is undefined.)
[…]
[This screen-shot shows an example of verb substitution suggestions that appear on the semantics diagram when the verb is unknown by the system.]This screen-shot shows an example of verb substitution suggestions that appear on the semantics diagram when the verb is unknown by the system. The semantics diagram has been created for ”frog,” ”hop,” ”to,” ”*lily” (meaning ”The frog hops to a lily.” A frog sketch is displayed in miniature under the word ”frog” to indicate that the object has been selected for the command. A question mark is displayed atop the diagram on the top left to indicate that there is an error in the diagram. Beneath ”hop” is written, ”unknown, try:” (meaning the word ”hop” is currently unknown to the system). Beneath is a tappable list of alternative verbs from which to choose, displayed as vertical red blocks. The verb suggestions are ”jump” and ”move.”
@@ Appendix D Abstract Semantic Structure
The following defines the basic structure of S2 (subsection 5.2), which is a simplified semantic structure graph representing an interpretable command. The concrete implementation of DrawTalking traverses this structure to generate final execution commands S3 (subsection 5.2).